OpenStudy (anonymous):

Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

x^2 + 3x + 1. = 0 Now as i was telling earlier, if we compare this equation with ax^2 + bx +c = 0, a = 1, b = 3 and c = 1. b^2-4ac = 9 - 4 = 5 Hence b^2-4ac is not a perfect square . (5 is not a perfect square.) b^2-4ac is used to determine the nature of roots. If it is not a perfect square, then the value of roots will not be integers and hence you cannot factorise the equation.